OpenStudy (anonymous):

3 questions left medal! please help

OpenStudy (anonymous):

7. Solve the absolute value equation. Graph the solution

OpenStudy (anonymous):



OpenStudy (anonymous):

3x - 5 = + or - 16/2

OpenStudy (anonymous):

c

OpenStudy (anonymous):

x = 13/3 x= 1

OpenStudy (anonymous):

A

OpenStudy (anonymous):

a. x= -1 5/6 or x= 5/6 b. x= -1 or x= 4 1/3 c. x= -1 5/6 or x= 4 1/3 d. x= -1 5/6 or x= -1

OpenStudy (anonymous):

@Loser66 do you think you could help me? :c

OpenStudy (loser66):

How to help?

OpenStudy (loser66):

Or, what kind of help you need? answer? walk you through?

OpenStudy (anonymous):

help step by step so I understand know how to work it? please

OpenStudy (loser66):

2|3x-5| -8 =8 right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

+8 both sides, get what?

OpenStudy (anonymous):

I get x= 13/3 and x= -1

OpenStudy (abb0t):

3 questions on what?

OpenStudy (loser66):

That's the very first step, I guide you step by step but you have to work on it for me to understand how far you are.

OpenStudy (anonymous):

algebra 2 I have 3 questions I need help with @abb0t

OpenStudy (anonymous):

@Loser66 oh okay

OpenStudy (anonymous):

well I used my calculator and that's what I got

OpenStudy (loser66):

As long as you know what you are doing, you are ok and the answers are good.

OpenStudy (anonymous):

oo okay! so the answer would be B right?

OpenStudy (loser66):

yup